OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x2 = −6x − 7?

OpenStudy (allank):

Hint: At the x-intercept(s), y=0. Plugging in that into the equation should get you the answer.

OpenStudy (anonymous):

Yes. All this is asking for is for you to find what x is when y = 0. \[y - x^2= -6x - 7 \implies 0 = x^2 - 6x - 7\]Can you factor it?

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

y − x^2 = −6x − 7 y = x^2 − 6x − 7 y = (x+1)(x-7) x+1=0 x=-1 x-7=0 x=7 so the answer is (−1, 0) and (7, 0)

OpenStudy (anonymous):

ok thats what i just got

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

@laxplayer Please don't just give the answer. This is a tutoring site and not give answer site. @simplybeyoutiful Good job :)

OpenStudy (anonymous):

@Calcmathlete look at this then Need help? Ask a question, and get an answer from a student just like you.

OpenStudy (anonymous):

Look at the code of conduct please: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

off from the website it self and @simplybeyoutiful great job:D

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

np :)